OpenStudy (mathmath333):

average salary for total employees is 2000.If the grade A employees's average salary is 4000, and grade B employee's average salary is 1250 then the total numbers of employees could be

OpenStudy (mathmath333):

a. 450 b. 300 c. 110 d. 500

ganeshie8 (ganeshie8):

suppose there are \(\large n\) total employees and out of which \(\large a\) employees are of grade \(\large a\) : \[\large 1250a + 4000(n-a) = 2000n\]

ganeshie8 (ganeshie8):

can you simplify that equation further ?

ganeshie8 (ganeshie8):

simplify it till the point you get this : \[\large 8n = 11a\]

ganeshie8 (ganeshie8):

since 11 is a prime and cannot divide 8, that means \(\large 11\) must divide \(\large n\) Look at your options - which option is a multiple of 11 ?

OpenStudy (mathmath333):

its c

ganeshie8 (ganeshie8):

Yep!

OpenStudy (mathmath333):

can i use digit sum trick here

ganeshie8 (ganeshie8):

you can.. all those tricks you had memorized are useful in working these problems quick

OpenStudy (mathmath333):

i found the ration 8a=3b

ganeshie8 (ganeshie8):

work it again

OpenStudy (mathmath333):

*ratio

ganeshie8 (ganeshie8):

but we're working a, n right ?

ganeshie8 (ganeshie8):

who is this b ?

OpenStudy (mathmath333):

b grade employees

ganeshie8 (ganeshie8):

why are u working b grade employees ? the equation we setup involves only a and n eh ?

OpenStudy (mathmath333):

ok

OpenStudy (mathmath333):

thanks very much

ganeshie8 (ganeshie8):

np :) you can simply replace a by n-b in ur equation : 8a = 3b

ganeshie8 (ganeshie8):

8(n-b) = 3b 8n = 11b

ganeshie8 (ganeshie8):

we get the same result again - n needs to be a multiple of 11

OpenStudy (mathmath333):

yes ur method makes sense